Using the string frequency equations when no tensi
Answer #1

For 2 Hz

=>   2 =   sqrt(T/u)/(2 * 1)

=>   sqrt(T/u) = 4

=> Tension = 4 N

=> linear mass density = 0.25 kg/m

For 3 Hz

=>   3 =   sqrt(T/u)/(2 * 1)

=>   sqrt(T/u) = 6

=> Tension = 6 N

=> linear mass density = 0.166 kg/m

For 4 Hz

=>   4 =   sqrt(T/u)/(2 * 1)

=>   sqrt(T/u) = 8

=> Tension = 8 N

=> linear mass density = 0.125 kg/m